### v1

@@ C.2 Miscellaneous Datasets
The Causal Chambers. Recently, Gamella et al. (2024) have proposed an interesting new type of dataset where the data is generated from real-world experiments but the ground-truth graph is known. Gamella et al. (2024) designed two computer-controlled physical simulators, that can generate observational and interventional data (i.i.d. as well as time series) with experimentally verified ground-truth graphs. Due to the recency of its development, as far as we know, no causal discovery method has been evaluated on Causal Chamber data yet (see (Lohse and Wahl, 2024) for an investigation on var-/$R^{2}$-sortability of Causal Chambers time series data).
@@ Appendix D Metrics to evaluate causal models
A major limitation of these metrics is that they are purely graphical without any notion of causality. Other structural metrics assess the distance in terms of topological ordering (e.g., Rolland et al. (2022)) conditional independencies (Textor et al., 2016), d-separation statements (Wahl and Runge, 2024), node-permutation tests (Eulig et al., 2023), etc. More focused on the effect of interventions, Peters and Bühlmann (2015) proposed the Structural Interventional Distance (SID) which counts the number of interventional distributions that would be wrongly computed using the parents from the learned structure as its adjustment sets. While considering interventions, this measure is still about the graph and correlates strongly with SHD (Gentzel et al., 2019). A generalization of this measure that considers other adjustment sets has also been proposed by Henckel et al. (2024). It also has the advantage of being directly applicable to CPDAGs as it returns a scalar instead of bounds and is computationally less demanding.
@@ Appendix E Causal Representation Learning
We conclude by stating that we only focused on causal representation learning, but the realm of domains where causal methods are applied has grown abundantly yielding many other possible applications. For example, causally-inspired algorithms have been proposed to tackle the problem of multidomain data (e.g., Arjovsky et al. (2019)), where different domains are interpreted as different interventional environments (for examples of datasets, see Gulrajani and Lopez-Paz (2020)). We could also mention the active field of causal reinforcement learning (e.g., Ahmed et al. (2020); Ke et al. (2021)).

### v2

@@ C.2 Miscellaneous Datasets
The Causal Chambers. Recently, Gamella et al. (2025) have proposed an interesting new type of dataset where the data is generated from real-world experiments but the ground-truth graph is known. Gamella et al. (2025) designed two computer-controlled physical simulators, that can generate observational and interventional data (i.i.d. as well as time series) with experimentally verified ground-truth graphs. Due to the recency of its development, as far as we know, no causal discovery method has been evaluated on Causal Chamber data yet (see Lohse and Wahl (2024) for an investigation on var-/$R^{2}$-sortability of Causal Chambers time series data).
@@
TableShift. Recently Gardner et al. (2024) proposed TableShift: a benchmark comprising many tabular datasets chosen to assess the robustness of machine learning methods to distribution shift. Although not originally intended for causal discovery, it has been explored in Nastl and Hardt (2024), where causal discovery methods are used with the aim of finding causal parents of variables of interest.
@@ Appendix D Metrics to evaluate causal models
A major limitation of these metrics is that they are purely based on the graph without any notion of causality. Other structural metrics assess the distance in terms of topological ordering (e.g., Rolland et al. (2022)) conditional independencies (Textor et al., 2016), d-separation statements (Wahl and Runge, 2025), node-permutation tests (Eulig et al., 2023), etc. More focused on the effect of interventions, Peters and Bühlmann (2015) proposed the Structural Interventional Distance (SID) which counts the number of interventional distributions that would be wrongly computed using the parents from the learned structure as its adjustment sets. While considering interventions, this measure is still about the graph and correlates strongly with SHD (Gentzel et al., 2019). A generalization of this measure that considers other adjustment sets has also been proposed by Henckel et al. (2024). It also has the advantage of being directly applicable to CPDAGs as it returns a scalar instead of bounds and is computationally less demanding.
@@ Appendix E Causal Representation Learning
We conclude by stating that we only focused on causal representation learning, but the realm of domains where causal methods are applied has grown abundantly yielding many other possible applications. While deviating from classical causal discovery, they often use more realistic datasets. Some new methods relying on LLM can directly rely on text metadata to learn relevant causal variables (Liu et al., 2024). Causally-inspired algorithms have also been proposed to tackle the problem of multidomain data (e.g., Arjovsky et al. (2019)), where different domains are interpreted as different interventional environments (for examples of datasets, see Gulrajani and Lopez-Paz (2021)). Finally, we could also mention the active field of causal reinforcement learning (e.g., Ahmed et al. (2021); Ke et al. (2021)).